OpenStudy (anonymous):

What is the distance between (-3, -4) and (5, 2) on the coordinate plane? 10 units 14 units 4 units 8 units

OpenStudy (anonymous):

\[\sqrt{\left( x2 - x1 \right)^{2} + \left( y2 - y1 \right)^{2}}\]

OpenStudy (goformit100):